The vertical and horizontal components of the initial velocity of 36m/s and 64m/s find the inclination to the horizontal at which the projectile was fired


1
Expert's answer
2021-01-31T18:57:55-0500
"\theta=\arctan{\frac{v_y}{v_x}}\\\theta=\arctan{\frac{64}{36}}=60.64\degree"